# Using matrix-product states for time-series machine learning

### v1

@@ Abstract
Matrix-product states (MPS) have proven to be a versatile ansatz for modeling quantum many-body physics. For many applications, and particularly in one-dimension, they capture relevant quantum correlations in many-body wavefunctions while remaining tractable (polynomial scaling) to store and manipulate on a classical computer. This has motivated researchers to also apply the MPS ansatz to machine learning (ML) problems where capturing complex correlations in datasets is also a key requirement. Here, for the first time, we develop and apply an MPS-based algorithm, MPSTime, for learning a joint probability distribution underlying an observed time-series dataset, and show how it can be used to tackle important time-series ML problems, including classification and imputation. MPSTime can efficiently learn complicated time-series probability distributions directly from data, requires only moderate maximum MPS bond dimension $\chi_{\rm{max}}$, with values for our applications ranging between $\chi_{\rm{max}}=20$-$150$, and can be trained for both classification and imputation tasks under a single logarithmic loss function. Using synthetic and publicly available real-world datasets—spanning applications in medicine, energy, and astronomy—we demonstrate performance competitive with state-of-the-art ML approaches, but with the key advantage of encoding the full joint probability distribution learned from the data. By sampling from the joint probability distribution and calculating its conditional entanglement entropy, we show how its underlying structure can be uncovered and interpreted. This manuscript is supplemented with the release of publicly available code package MPSTime that implements our approach, and can be used to reproduce all presented results. The efficiency of the MPS-based ansatz for learning complex correlation structures from time-series data is likely to underpin interpretable advances to a broad range of challenging time-series ML problems across science, industry, and medicine.
@@ I Introduction
The ability of MPS to express a range of correlation functions while remaining tractable for computation makes it an ideal foundation for developing machine-learning (ML) algorithms. Here we use the term machine learning to refer to a class of statistical algorithms for tackling data-driven problems, like classification, that involves learning structures from data. The use of MPS for ML was realized by Stoudenmire and Schwab [15], who showed that MPS can be used for classifying numerical digits from (two-dimensional) images of human handwriting. More specifically, the wavefunction of a spin-$1/2$ chain stored using an MPS can be used to classify handwritten digits in the MNIST dataset with an error $<1\%$. In this approach real-valued data is encoded into an exponentially large Hilbert space using a nonlinear encoding function. A classifier is then trained by finding an MPS for each class having large overlap with data from that class while being orthogonal with other classes. This approach was subsequently extended to perform unsupervised generative ML with the MNIST handwriting image classification dataset [16, 17, 18, 19], among others [20, 21].
@@
More recently, MPS-based generative models have been applied to tackle a diverse range of important unsupervised ML problems such as anomaly detection [22, 23, 24], image segmentation [25], and clustering [26], and others [27, 28]. Beyond conventional ML applications, MPS-based generative models have also been applied to the problem of compressed sensing [19], where they have demonstrated strong performance in sparse signal reconstruction. In the context of supervised ML, MPS-based classifiers have been applied to a multitude of real-world datasets including weather data [29], medical images [25], audio [30], and particle event detection [31, 32].
@@
A key advantage of using MPS over deep-learning architectures for ML is their inherent interpretability [24]. With MPS, the learned correlations are explicitly encoded in the model’s tensors as conditional probabilities, allowing one to directly extract meaningful insights into the model’s behavior and its representation of dependencies within a dataset. For example, analyzing the entanglement properties of an MPS trained on image data allowed researchers to identify which features were most crucial for identifying samples [27]. In unsupervised anomaly detection, MPS achieved performance competitive with traditional deep-learning approaches such as variational autoencoders (VAEs) and generative adversarial networks (GANs), while offering considerably richer explainability [24]. By computing interpretable properties of the MPS such as the von Neumann entropy and mutual information, researchers uncovered the informational content of individual features and their contextual dependencies, shedding light on which features were most important for identifying anomalous patterns. This transparency provides a stark contrast to the opaque nature of deep learning architectures, which, while highly expressive, are challenging to understand and interpret [33].
@@
The quantification of complex correlations in time series is central to data-driven time-series modeling, which underpins a range of time-series ML tasks, from diagnosing disease from measured brain signals, to forecasting inflation rates from economic data. There are many existing algorithmic approaches for tackling time-series analysis tasks, including time-series modeling, classification, anomaly detection, imputation, and forecasting [34]. These methods range from traditional statistical methods, such as autoregressive (AR) models, which are tractable but can capture only relatively simple temporal structures, to highly flexible and powerful so-called ‘black-box’ approaches based on deep neural networks, which can capture much more complex and long-range temporal structures (but are challenging to interpret). Univariate time series, considered here to be uniformly sampled (at a constant sampling period $\Delta t$) and thus representable as the vector $x_{t}$ (for $t=1,\dots,T$), encode potentially complex one-dimensional (temporal) correlation structures, similar to the one-dimensional (spatial) correlation structures of wavefunctions in one-dimensional quantum systems that can be approximated using MPS. The success of MPS in approximating wavefunctions in one-dimensional quantum systems and recent high-performing ML applications motivates us to extend this promising and flexible algorithmic framework to tackle time-series ML problems. We note that the one-dimensional nature of the dataset does not guarantee the applicability of MPS. For example, text data, which is inherently one dimensional has volume-law mutual information growth [35] and therefore cannot be efficiently modeled using MPS. Nevertheless, time-series data arise both in nature and in technological applications from a diverse array of sources. This means that the correlation structure of the data is also likely to vary significantly. In this paper, we develop and apply an MPS-based algorithm for learning a joint probability distribution underlying a temporal process directly from an observed time-series dataset. We show that the MPS ansatz is indeed an efficient and powerful one for capturing a diverse range of complex temporal correlation structures, and further show how it can be used as the basis for novel algorithms that we introduce for tackling time-series ML problems, focusing on time-series classification (the inference of a categorical label from a time series) and imputation (the inference of a subset of unobserved values from a time series). We implement this algorithm and provide our classification and imputation tools in the publicly available software package MPSTime [^1] https://github.com/jmoo2880/MPSTime.jl.
@@
Extending the existing MPS-based ML framework of Stoudenmire and Schwab [15], which was applied to two-dimensional images, to time series requires substantial further development. While MNIST figures are characterized by grayscale pixels with approximately continuous grayscale values, significant progress can be made by approximating the pixels as either black or white [16]. On the other hand, for time-series data, the real-valued amplitudes, $x_{t}$, are crucial to properly represent in order to appropriately encode the statistical structures of the underlying dynamical process. This means that a carefully chosen encoding process is required to map the real-valued time-series data into finite dimensional vectors that can be interfaced with MPS. In fact, it was recently shown that MPS can approximate any probability density function for continuous real-valued data [20]. A second key aspect of our algorithm is that we can use the same loss function for training the MPS both for classification and generative ML. This means that our method can be used to learn a probabilistic model that encodes the joint probability distribution of a data class, and also train it to act as a classifier. This unified framework for time-series ML stands in contrast to the existing algorithmic literature on time-series analysis that, due to the typical intractability of learning high-dimensional joint distributions from data, is highly disjoint in its development of different algorithms for different classes of statistical time-series problems.
@@
This paper is structured as follows. In Sec. II we cover the relevant MPS theory and specify the tensor network structure we used in subsequent investigations. In Sec. II.1 we present an approach for encoding real-valued time-series data into a finite-dimensional Hilbert space with nonlinear feature maps. Section II.3 and Sec. II.4 describe the methods we use to perform time-series imputation and classification, respectively. In Sec. III, we firstly synthesize datasets to demonstrate the generative modeling and imputation capabilities of our MPS algorithm, MPSTime. We then introduce three real-world datasets focusing on key application areas: energy, medicine and physics, to showcase the ability of MPSTime to perform both classification and imputation of missing data points. Moving further, in Sec. IV we highlight the interpretability of the joint probability distribution encoded in the trained MPS via sampling trajectories and by calculating conditional entanglement entropies. Finally, Sec. V contains the discussion and conclusion.
II The Matrix Product State Formalism
@@
In quantum many-body physics, a system of $T$ particles can be modeled by a wavefunction $\Psi(s_{1},s_{2},\ldots,s_{T})$, which assigns a probability amplitude to each possible configuration of the many-body system. For any specific configuration $\mathbf{s}=(s_{1},s_{2},\ldots,s_{T})$, Born’s rule states that the squared-norm of the wavefunction $p(\mathbf{x})=\lvert\Psi(\mathbf{s})\rvert^{2}$ defines the probability density of observing the configuration. Here, the wavefunction encapsulates the full joint distribution over an exponentially large space of $d^{T}$ possible states, where $d$ is the number of local states that each particle can take (e.g., $d=2$ for spin-1/2 particles).
@@
Figure 1: Mapping between quantitative formulations of quantum many-body physics and time-series analysis. Concepts in quantum physics (left, blue shading) and time-series analysis (right, red shading), and their associated theoretical and analytic tools, share similarities that motivate our MPS-based approach to time-series ML. (a) A $d$-level spin chain (left) exhibits one-dimensional (1D) spatial ordering, analogous to the 1D temporal ordering of time series, in which each sample is mapped to a discrete state which parallels that of an individual $d$-level quantum spin. (b) The probability density captured by the square of the wavefunction $|\Psi(\mathbf{x})|^{2}$ (left), is analogous to the joint probability density of a time-series process $p(\mathbf{x})$ (right). The wavefunction is a vector in an exponentially large Hilbert space, while time-series probability density is a distribution in the space of possible time series. A single measurement outcome (i.e., one spin chain configuration $\mathbf{x}$) governed by Born’s rule corresponds to sampling a single time-series realization (i.e., one time-series instance $\mathbf{x}$) from its generative distribution. (c) The entanglement entropy (left), which quantifies the degree of quantum entanglement between two spatial subsystems, is conceptually related to classical quantities such as the mutual information between temporal segments (right), both capturing statistical dependencies in their respective domains.
@@
As we will show, for the purposes of time-series analysis, it is advantageous to view a $T$-sample time-series instance $\mathbf{x}=(x_{1},x_{2},\dots,x_{T})$ as a specific configuration of a $T$-body system, where each temporal observation $x_{t}$ – originally real-valued and continuous – is mapped to a discrete state $s_{t}$. Throughout the text, we refer to a single temporal observation $x_{t}$ as a time-series amplitude or time-series value. Here, analogous to a quantum many-body system, the wavefunction $\Psi(x_{1},x_{2},\dots,x_{T})$ encodes the joint distribution of time-series data, and is determined by the properties of the underlying generative process. Analogous ideas in quantum many-body physics and time-series machine learning are illustrated in Fig. 1. We emphasize that the joint probability distribution encompasses not just the empirical distribution of the observed time-series, but the full generative distribution from which individual time-series instances are sampled. In other words, each time-series instance $\mathbf{x}$ of $T$ measurements serves as a point sample from a $T$-dimensional distribution. As with its quantum analogue, within our MPS formalism for ML the probability density of observing a time-series instance $\mathbf{x}$ from the distribution encoded by $\Psi$ is given by Born’s rule:
@@
Although such a wavefunction is exponentially large in general, it may be approximated by a $T$-site MPS ansatz [4] which provides a compressed tensor network representation:
@@
Here, $W$ is the low-rank MPS approximation of the original wavefunction (which can also be expressed in tensor form), each site $A^{s_{t}}$ corresponds to a measured point in time $x_{t}$, and the dimension of the bond indices $\bm{\alpha}=\{\alpha_{1},\ldots,\alpha_{T-1}\}$ may be adjusted to tune the maximum complexity of the MPS ansatz. Crucially, by compressing the wavefunction, and by extension, the joint distribution it encodes, the MPS approximation offers an explicit and tractable model of the generative distribution underlying time-series data. The ability to work directly with a tractable model of the full time-series joint distribution – rather than relying solely on the empirical distribution of observations – is what underpins the key innovations of our present work on MPS-based algorithms for time-series analysis. In particular, we show how inferring a joint distribution of time series with MPS, analogous to the distribution encoded by the quantum many-body wavefunction, allows us to tackle important statistical learning problem classes, including classification and imputation.
@@
Figure 2: Time-series machine learning with MPS. (a) Each Real-valued time series amplitude $x_{t}$ is encoded in a $d$-dimensional vector $\phi_{t}$ by projecting its value onto a truncated orthonormal basis with $d$ basis functions. An entire time series (of length $T$ samples) is then encoded as a set of $T$ $\phi_{t}$ vectors, which we represent as a product state embedded in a $d^{T}$ dimensional Hilbert space. (b) Using observed time-series from a dataset, a generally entangled MPS – depicted here using Penrose graphical notation – with maximum bond dimension $\chi_{\rm{max}}$ is trained with a DMRG-inspired sweeping optimization algorithm to approximate the underlying joint distribution of the data. Two copies of the trained MPS (one conjugate-transposed, denoted by the dagger $\dagger$) with open physical indices encodes the learned distribution, allowing us to sample from and do inference with complex high-dimensional distributions in a tractable manner. In this work, we introduce MPS-based learning algorithms for two important time-series ML problems: (c) imputation (inferring unmeasured values of a time series), and (d) classification (inferring a time-series class). (c) Generative time-series modeling: we use conditional sampling to perform imputation of missing datapoints. Known points of a time series (black lines) project the MPS into a subspace, which is then used to find the unknown datapoints (red line). The same method can be used to tackle some forecasting problems if the missing points are future values. (d) MPS for classification: multiple labeled classes of time series are used to train MPSs. Taking the overlap of unlabeled time-series data (encoded as a product state) with each MPS determines its class.

### v2

@@ Abstract
Matrix-product states (MPS) have proven to be a versatile ansatz for modeling quantum many-body physics. For many applications, and particularly in one-dimension, they capture relevant quantum correlations in many-body wavefunctions while remaining tractable (polynomial scaling) to store and manipulate on a classical computer. This has motivated researchers to also apply the MPS ansatz to machine learning (ML) problems where capturing complex correlations in datasets is also a key requirement. Here, for the first time, we develop an MPS-based algorithm, MPSTime, for learning a joint probability distribution underlying a time-series dataset, and show how it can be used to tackle important time-series ML problems, including classification and imputation. MPSTime can efficiently learn complicated time-series probability distributions directly from data, requires only moderate maximum MPS bond dimension $\chi_{\rm{max}}$ (with values for our applications ranging between $\chi_{\rm{max}}=20-160$), and can be trained for both classification and imputation tasks under a single logarithmic loss function. Using synthetic and publicly available real-world datasets—spanning applications in medicine, energy, and astronomy—we demonstrate performance competitive with state-of-the-art ML approaches, but with the key advantage of encoding the full joint probability distribution learned from the data, which is useful for analyzing and interpreting its underlying structure. This manuscript is supplemented with the release of a publicly available code package MPSTime that implements our approach, and can be used to reproduce all presented results. The effectiveness of the MPS-based ansatz for capturing complex correlation structures in time-series data makes it a powerful foundation for tackling challenging time-series analysis problems across science, industry, and medicine.
@@ I Introduction
The ability of MPS to express a range of correlation functions while remaining computationally tractable makes it an ideal foundation for developing machine-learning (ML) algorithms. We use the term ‘machine learning’ to describe a class of statistical algorithms designed to solve data-driven problems, such as classification, by learning generalizable patterns and structures from data. The use of MPS for ML was first realized by Stoudenmire and Schwab [15], who showed that MPS can be used to classify numerical digits from (two-dimensional) images of human handwriting. More specifically, the wavefunction of a spin-$1/2$ chain stored using an MPS can be used to classify handwritten digits in the MNIST dataset with an error $<1\%$. In this approach real-valued data is encoded into an exponentially large Hilbert space using a nonlinear encoding function. A classifier is then trained by finding an MPS for each class (i.e., digit) having large overlap with data from that class, while being orthogonal with other classes. This approach was subsequently extended to perform unsupervised generative ML with the MNIST handwriting image classification dataset [16, 17, 18, 19], and others [20, 21].
@@
More recently, MPS-based generative models have been applied to tackle a diverse range of important unsupervised ML problems such as anomaly detection [22, 23, 24], image segmentation [25], and clustering [26], among others [27, 28]. Beyond conventional ML applications, MPS-based generative models have also been applied to the problem of compressed sensing [19], where they have demonstrated strong performance in sparse signal reconstruction. In the context of supervised ML, MPS-based classifiers have been applied to a multitude of real-world datasets including weather data [29], medical images [25], audio [30], and particle events [31, 32], highlighting their utility across a range of problems, and for various scientific disciplines.
@@
A key advantage of using MPS over conventional deep-learning models for ML is their inherent interpretability [24]. With MPS, the learned correlations are explicitly encoded in the model’s tensors as conditional probabilities, allowing one to directly extract meaningful insights into the model’s behavior and its representation of dependencies within a dataset. For example, analyzing the entanglement properties of an MPS trained on MNIST image data allowed researchers to identify which features were most crucial for identifying samples [27]. In unsupervised anomaly detection, MPS achieved performance competitive with flexible deep-learning data modeling algorithms, such as variational autoencoders (VAEs) and generative adversarial networks (GANs), while offering considerably richer explainability [24]. By computing interpretable properties of the MPS such as the von Neumann entropy and mutual information, researchers uncovered the informational content of individual features and their contextual dependencies, shedding light on which features were most important for identifying anomalous patterns. This transparency provides a stark contrast to the opaque nature of deep-learning architectures, which, while highly expressive, are often challenging to understand and interpret [33].
@@
The quantification of complex correlations in time series is central to data-driven time-series modeling, which underpins a range of time-series ML tasks, from diagnosing disease from measured brain signals, to forecasting inflation rates from economic data. There are many existing algorithmic approaches for tackling time-series analysis tasks, including time-series modeling, classification, anomaly detection, imputation, and forecasting [34]. These methods range from traditional statistical methods, such as autoregressive (AR) models, which are tractable but can capture only relatively simple temporal structures, to highly flexible and powerful so-called ‘black-box’ approaches based on deep neural networks, which can capture much more complex and long-range temporal structures (but are challenging to interpret). The latter approach encompasses a range of powerful and flexible time-series ML algorithms, including transformers [35], variational autoencoders (VAEs) [36], generative adversarial networks (GANs) [37], and denoising diffusion probabilistic models [38]. Univariate time series, considered here to be uniformly sampled (at a constant sampling period $\Delta t$), and thus representable as the vector $x_{t}$ (for $t=1,\dots,T$), encode potentially complex one-dimensional (temporal) correlation structures, similar to the one-dimensional (spatial) correlation structures of wavefunctions in one-dimensional quantum systems that can be approximated using MPS. The success of MPS in approximating wavefunctions in one-dimensional quantum systems and recent high-performing ML applications motivates us to extend this promising and flexible algorithmic framework to tackle time-series ML problems. In this paper, we develop and apply an MPS-based algorithm, MPSTime, for learning a joint probability distribution directly from an observed time-series dataset. We show that the MPS ansatz is an efficient and powerful one for capturing a diverse range of complex temporal correlation structures, and further show how it can be used as the basis for novel algorithms that we introduce for tackling time-series ML problems, focusing here on time-series classification (the inference of a categorical label from a time series) and imputation (the inference of a subset of unobserved values from a time series). We implement this algorithm and provide our classification and imputation tools in the publicly available software package MPSTime [^1] https://github.com/hugopstackhouse/MPSTime.jl.
@@
Extending the existing MPS-based ML framework of Stoudenmire and Schwab [15] – which was applied to two-dimensional images – to univariate time series requires substantial further development. While MNIST figures are characterized by grayscale pixels with approximately continuous grayscale values, significant progress can be made by approximating the pixels as either black or white [16]. On the other hand, for time-series data, it is crucial to properly represent the real-valued amplitudes, $x_{t}$, in order to appropriately encode the statistical structure contained in the dataset. This means that a carefully chosen encoding process is required to map each real-valued time-series datapoint into a finite dimensional vector that can be interfaced with MPS. In fact, using an appropriate encoding, it was recently shown that the MPS ansatz can approximate any probability density function for continuous real-valued data [20]. A second key aspect of our MPSTime algorithm is that we can use the same loss function for training the MPS both for classification and generative ML (e.g., synthetic time-series generation, forecasting, or imputation). This means that MPSTime can be used to infer a probabilistic model of the joint probability distribution of a time-series data class, which can be used as the basis of a time-series classification model. This unified framework for time-series ML stands in contrast to the existing algorithmic literature on time-series analysis that, due to the typical intractability of learning high-dimensional joint distributions from data, is highly disjoint in its development of different algorithms for different classes of statistical time-series problems.
@@
This paper is structured as follows. In Sec. II we cover the relevant MPS theory and specify the tensor network structure we used in subsequent investigations. In Sec. II.1 we present an approach for encoding real-valued time-series data into a finite-dimensional Hilbert space with nonlinear feature maps. Section II.3 and Sec. II.4 describe the methods we use to perform time-series imputation and classification, respectively. In Sec. III, we firstly synthesize datasets to demonstrate the generative modeling and imputation capabilities of MPSTime. We then introduce three real-world datasets focusing on key application areas –energy, medicine, and physics – to showcase the ability of MPSTime to perform both classification and imputation of missing data points. Moving further, in Sec. IV we highlight the interpretability of the joint probability distribution encoded in a trained MPS by sampling trajectories from the trained model and calculating conditional entanglement entropies. By performing imputation on out-of-distribution instances, we also shed light on the representation of the data distribution learned by the MPS. Broader implications of the work are discussed in Sec. V.
II The Matrix-Product State Formalism
@@
In quantum many-body physics, a system of $T$ particles can be modeled by a wavefunction $\Psi(s_{1},s_{2},\ldots,s_{T})$, which assigns a probability amplitude to each possible configuration of the many-body system. For any specific configuration $\mathbf{s}=(s_{1},s_{2},\ldots,s_{T})$, Born’s rule states that the squared-norm of the wavefunction i.e., $p(\mathbf{s})=\lvert\Psi(\mathbf{s})\rvert^{2}$ defines the probability density of observing the configuration. Here, the wavefunction encapsulates the full joint distribution over an exponentially large space of $d^{T}$ possible states, where $d$ is the number of local states that each particle can take (e.g., $d=2$ for spin-1/2 particles).
@@
Figure 1: Mapping between quantitative formulations of quantum many-body physics and time-series analysis. Concepts in quantum physics (left, blue shading) and time-series analysis (right, red shading), and their associated theoretical and analytic tools, share similarities that motivate our MPS-based approach to time-series ML. (a) A $d$-level spin chain (left) exhibits one-dimensional (1D) spatial ordering, analogous to the 1D temporal ordering of time series (right). In this work, we propose mapping each time-series sample to a discrete state, similar to an individual $d$-level quantum spin. (b) The probability density captured by the square of the wavefunction $|\Psi(\mathbf{s})|^{2}$ (left), is analogous to the joint probability density $p(\mathbf{x})$ (right), which assigns a probability to every possible time series. Here, the square of the wavefunction encodes a probability distribution over the space of possible measurement outcomes, while the joint density $p(x)$ defines a distribution over the space of possible time series (i.e., $\mathbb{R}^{T}$). A single measurement outcome (i.e., one spin chain configuration $\mathbf{s}$) governed by Born’s rule corresponds to sampling a single time-series realization (i.e., one time-series instance $\mathbf{x}$) from its generative distribution. (c) The entanglement entropy (left), which quantifies the degree of quantum entanglement between two spatial subsystems (e.g., $A$ and $B$), is conceptually related to classical quantities such as the mutual information (right) between temporal segments (e.g., $A$ and $B$), both capturing statistical dependencies in their respective domains.
@@
As we will show, for the purposes of time-series analysis, it is advantageous to view a $T$-sample time-series instance $\mathbf{x}=(x_{1},x_{2},\dots,x_{T})$ as a specific configuration of a $T$-body system, as in Fig. 1(a), where each temporal observation $x_{t}$ – originally real-valued and continuous – is mapped to a discrete state $s_{t}$. Throughout the text, we refer to a single temporal observation $x_{t}$ as a time-series amplitude or time-series value. Here, analogous to a quantum many-body system, the data wavefunction $\Psi(x_{1},x_{2},\dots,x_{T})$ encodes the joint distribution of time-series data, and is determined by the properties of the underlying generative process. Analogous ideas in quantum many-body physics and time-series machine learning are illustrated in Fig. 1. In this probabilistic picture, each time-series instance $\mathbf{x}$ of $T$ measurements thereby serves as a point sample from a $T$-dimensional joint distribution, as depicted in Fig. 1(b). As with its quantum analogue, within our MPS formalism for ML the probability density of observing a time-series instance $\mathbf{x}$ from the distribution encoded by $\Psi$ is given by Born’s rule:
@@
Although such a wavefunction is exponentially large in general, it may be approximated by a $T$-site MPS ansatz [4] which provides a compressed representation:
@@
Here, $W$ is the low-rank MPS approximation of the original wavefunction (which can also be expressed in tensor form), each site $A^{s_{t}}$ corresponds to a measured point in time $x_{t}$, and the dimension of the bond indices $\bm{\alpha}=\{\alpha_{1},\ldots,\alpha_{T-1}\}$ may be adjusted to tune the maximum complexity of the MPS ansatz. Crucially, by compressing the wavefunction, and by extension, the joint distribution it encodes, the MPS approximation offers an explicit and tractable model of the generative distribution underlying observed time-series data. Working directly with an approximation of the time-series joint distribution – rather than relying solely on the empirical distribution of observations – is what sets our MPS-based approach apart from existing methods. In particular, we show how inferring a joint distribution of time series with MPS, analogous to the joint distribution of states encoded by the quantum many-body wavefunction, allows us to tackle important statistical learning problem classes, including classification and imputation. Further, as in Fig. 1(c), by drawing on the breadth of existing tools for analyzing MPS in quantum many-body physics, we show that interpretable quantities that would otherwise be highly challenging or impossible to obtain from the empirical dataset alone, can be straightforwardly extracted from the MPS.
@@
Figure 2: MPSTime, a framework for time-series machine learning with Matrix-Product States (MPS). (a) Encoding: Each real-valued time series amplitude $x_{t}$ is encoded in a $d$-dimensional vector $\phi_{t}$ by projecting its value onto a truncated orthonormal basis with $d$ basis functions. An entire time series (of length $T$ samples) is then encoded as a set of $T$ $\phi_{t}$ vectors, which we represent as a product state embedded in a $d^{T}$ dimensional Hilbert space. (b) MPS training: Using observed time series from a dataset, a generally entangled MPS – depicted here using Penrose graphical notation – with maximum bond dimension $\chi_{\rm{max}}$ is trained with a DMRG-inspired sweeping optimization algorithm to approximate the joint distribution of the training data. Two copies of the trained MPS (one conjugate-transposed, denoted by the dagger $\dagger$) with open physical indices encodes the learned distribution, allowing us to sample from and do inference with complex high-dimensional time-series distributions. In this work, we introduce MPS-based learning algorithms, which we collectively refer to as MPSTime, for two important time-series ML problems: (c) imputation (inferring unmeasured values of a time series), and (d) classification (inferring a time-series class). (c) Generative time-series modeling: we use conditional sampling to perform imputation of missing datapoints. Known points of a time series (black lines) project the MPS into a subspace, which is then used to find the unknown datapoints (red line). The same method can be used to tackle some forecasting problems if the missing points are future values. (d) MPS for classification: multiple labeled classes of time series are used to train MPSs. Taking the overlap of unlabeled time-series data (encoded as a product state) with each MPS determines its class.